OpenStudy (anonymous):

Plz help what is the solutions tu this equation x^2+4x=-2

OpenStudy (mathteacher1729):

"Please help. What is the solution to this equation?" \[x^2+4x=-2\] We can do this a few ways. We can graph it and find where the curve \(y=x^2+4x\) intersects the line \(y =-2\). (see attached image). Or we can graph \(y = x^2+4x+2\) and see where this intersects the x-axis (x-axis is the line y = 0). (see other attached picture). You can also use factoring, but that takes a bit of time to explain. Let me know if the graphs make sense. :)

OpenStudy (anonymous):

Thanxx